OpenStudy (anonymous):

I would appreciate any help with this problem "Find the equation of the tangent line to the graph of f at the given point: f(t) = (t-4)(t^2-6), at (2,4)

OpenStudy (anonymous):

I think I start this problem by doing the product rule, but I'm not to sure

OpenStudy (turingtest):

yes, to find the tangent at a point you will need the derivative at that point, so find f'(t) and yes, it will require the product rule

OpenStudy (anonymous):

Once I finish doing the product rule, would I just plug in the point ?

OpenStudy (anonymous):

Couldn't that expression be simplified into a single polynomial?

OpenStudy (anonymous):

"would I just plug in the point ?" - Yes, but only the t=2 part; you can ignore the f(2)=4 for now.

OpenStudy (anonymous):

Okay and when I find that answer would I be finished ?

OpenStudy (anonymous):

or do I plug that into y=mx+b?

OpenStudy (anonymous):

Not quite. That will get you the slope of the line, but you still need an equation for the line.

OpenStudy (anonymous):

Yes, you'll have the slope and a point, so you can either use y=mx+b or y-y1=m(x-x1)

OpenStudy (anonymous):

So I used y=mx+b and I got a final answer of y=-10x+24. Would you mind checking my answer ?

OpenStudy (anonymous):

Alright, let me get my pencil . . .

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

Ok, slope is correct...

OpenStudy (anonymous):

Yes, that is the tangent line at that point. Good job!

OpenStudy (anonymous):

Did you use the product rule?

OpenStudy (anonymous):

woohoo! thank you very much for your help (:

OpenStudy (anonymous):

In most cases when you have a product, product rule is the way to go, but this one easily simplified to t^3 -4t^2 -6t +24, so you could also differentiate term by term.